# LaM-SLidE: Latent Space Modeling of Spatial Dynamical Systems via Linked Entities

### v3

@@ Appendix F Proofs
Proof. Assume $|E|>|{\mathcal{I}}|$. By pigeonhole principle, any function $f:E\mapsto{\mathcal{I}}$ must map at least two distinct elements of $E$ to the same element in $I$, Therefore $f$ cannot be injective.
@@
Therefore, an injective identifier assignment function $\mathtt{ida}(\cdot)\in I$ only exists if $|E|\leqslant|{\mathcal{I}}|$. Hence the set I is non-empty in this case and empty otherwise.
F.2 Proof of Proposition 3.4
@@
Where $(m)_{n}$ is commonly referred as falling factorials, the number of injective functions from a set of size $n$ to a set of size $m$.
@@ G.1 Number of Parameter
We conducted scaling experiments on both the MD17 and the Tetrapeptides (4AA) datasets to evaluate how LaM-SLidE ’s performance scales with model size. On MD17, we evaluate LaM-SLidE using model variants with 1.7M, 2.1M, and 2.5M parameters. Our results show that, for nearly all molecules, performance improves with parameter count in terms of ADE/FDE, see Table 13. Similarly, on the Tetrapeptides dataset, we evaluate using model variants with 4M, 7M, 11M, and 28M parameters. All performance metrics show consistent improvement with increased model capacity, see Table 14. These findings indicate favorable scaling behavior of our method.
@@
Table 14: Method comparison for predicting MD trajectories of tetrapeptides. The columns denote the JSD between distributions of torsion angles (backbone (BB), side-chain (SC) and all angles), the TICA, the MSM metric, and the number of parameters.
@@ G.2 Number of Function Evaluations (NFEs)
We conduct a comparative analysis on computational efficiency by measuring the number of function evaluations (NFEs) required to achieve the performance results reported in the main section of our publication. As shown in Table 15, our approach demonstrates remarkable efficiency compared to the previous state-of-the-art method, GeoTDM [35], across N-Body simulations, MD17 molecular dynamics, and ETH pedestrian motion forecasting experiments. Our model consistently requires significantly fewer NFEs than GeoTDM to reach comparable or superior performance levels.
It is worth noting that flow-based models generally require fewer NFEs compared to diffusion-based approaches like GeoTDM. However, this efficiency advantage does not come at the expense of performance quality [27]. Indeed, the relationship between NFEs and performance is not strictly monotonic, as demonstrated in other domains. For instance, Esser et al. [27] achieved optimal image generation results in terms of FID [36] with 25 NFEs, showing that computational efficiency and high performance can be simultaneously achieved with properly designed architectures.
Further, in the case of the Tetrapeptides (4AA) experiments shown in Section 4.5, we use an adaptive step size solver to reach the reported performance, which achieved better results than a Euler solver. We use Dopri5 as implemented in the torchdiffeq package [20].
Table 15: Comparison of the number of functions evaluations (NFEs) for LaM-SLidE and GeoTDM.
@@ G.3 Number of Learned Latent Vectors
We conducted experiments to quantify the relationship between model performance and the number of latent vectors $L$ using the MD17 dataset. As shown in Table 16, performance increases with the number of latent vectors $L$. Of particular significance is the performance at $L=21$, which corresponds to the maximum number of entities, allowing us to investigate whether this constitutes an upper bound on model capacity. Notably, performance continues to improve at $L=32$, indicating that model capacity scales favorably even beyond the number of entities. Still at $L=16$, representing a compressed latent representation, our model remains competitive with the second best method GeoTDM [35].
We further analyze if the improvement by increasing $L$ is due to the reconstruction performance of the encoder-decoder only. Figure 8 shows the reconstruction error for varying number of latent vectors $L$. Even with substantially fewer latent vectors than entities, the model achieves good reconstruction performance. This gap suggests, that the performance gains from increasing $L$ are not due to improved reconstruction, but from the ability of the model to leverage the enlarged latent space representation better.
@@ G.4 Identifier Pool Size
We evaluate the impact of identifier pool size using the MD17 dataset. This dataset contains at most 21 atoms, so in general an identifier pool of $|{\mathcal{I}}|=21$ would be enough. However, for the results shown in the main paper, we used $|{\mathcal{I}}|=32$. To investigate the impact of a larger identifier pool ${\mathcal{I}}$, we conduct additional experiments by training multiple first-stage models with varying identifier pool sizes and report the reconstruction error measured by Euclidean distance. Figure 7 shows that the reconstruction error increases with the size of the identifier pool, since a larger pool results in fewer updates to each entity embedding during training.
@@ G.5 Identifier Assignment
To assess the impact of different ID assignments on reconstruction performance, we run our model five times with different random ID assignments and measure the standard deviation across these assignments in terms of reconstruction error in Å. Results are shown in Figures 7 and 8. The low standard deviation across different ID assignments demonstrates the robustness of our model with respect to random ID assignment.

### v4

@@ Appendix F Proofs
Proof. Assume $|E|>|{\mathcal{I}}|$. By pigeonhole principle, any function $f:E\mapsto{\mathcal{I}}$ must map at least two distinct elements of $E$ to the same element in $I$. Therefore, $f$ cannot be injective.
@@
Therefore, an injective identifier assignment function $\mathtt{ida}(\cdot)\in I$ only exists if $|E|\leqslant|{\mathcal{I}}|$. Hence, the set I is non-empty in this case and empty otherwise.
F.2 Proof of Proposition˜3.4
@@
Where $(m)_{n}$ is commonly referred to as falling factorials, the number of injective functions from a set of size $n$ to a set of size $m$.
@@ G.1 Number of Parameter
We conducted scaling experiments on both the MD17 and the Tetrapeptides (4AA) datasets to evaluate how LaM-SLidE ’s performance scales with model size. On MD17, we evaluate LaM-SLidE using model variants with 1.7M, 2.1M, and 2.5M parameters. Our results show that, for nearly all molecules, performance improves with parameter count in terms of ADE/FDE, see Table˜13. Similarly, on the Tetrapeptides dataset, we evaluate using model variants with 4M, 7M, 11M, and 28M parameters. All performance metrics show consistent improvement with increased model capacity, see Table˜14. These findings indicate the favorable scaling behavior of our method.
@@
Table 14: Method comparison for predicting MD trajectories of tetrapeptides. The columns denote the JSD between distributions of torsion angles (backbone ($B$), side-chain (SC), and all angles), the TICA, the MSM metric, and the number of parameters.
@@ G.2 Number of Function Evaluations (NFEs)
We conduct a comparative analysis on computational efficiency by measuring the number of function evaluations (NFEs) required to achieve the performance results reported in the main section of our publication. As shown in Table˜15, our approach demonstrates remarkable efficiency compared to the previous state-of-the-art method, GeoTDM [35], across N-Body simulations, MD17 molecular dynamics, and ETH pedestrian motion forecasting experiments. Our model consistently requires significantly fewer NFEs than GeoTDM to reach comparable or superior performance levels.
It is worth noting that flow-based models generally require fewer NFEs compared to diffusion-based approaches, such as GeoTDM. However, this efficiency advantage does not come at the expense of performance quality [27]. Indeed, the relationship between NFEs and performance is not strictly monotonic, as demonstrated in other domains. For instance, Esser et al. [27] achieved optimal image generation results in terms of FID [36] with 25 NFEs, demonstrating that computational efficiency and high performance can be achieved simultaneously with properly designed architectures.
Furthermore, in the case of the Tetrapeptides (4AA) experiments shown in Section˜4.5, we employ an adaptive step size solver to achieve the reported performance, which yields better results than an Euler solver. We use Dopri5 as implemented in the torchdiffeq package [20].
Table 15: Comparison of the number of function evaluations (NFEs) for LaM-SLidE and GeoTDM.
@@ G.3 Number of Learned Latent Vectors
We conducted experiments to quantify the relationship between model performance and the number of latent vectors $L$ using the MD17 dataset. As shown in Table˜16, performance increases with the number of latent vectors $L$. Of particular significance is the performance at $L=21$, which corresponds to the maximum number of entities, allowing us to investigate whether this constitutes an upper bound on model capacity. Notably, performance continues to improve at $L=32$, indicating that model capacity scales favorably even beyond the number of entities. At $L=16$, representing a compressed latent representation, our model remains competitive with the second-best method, GeoTDM [35].
We further analyze whether the improvement resulting from increasing $L$ is due solely to the encoder-decoder’s reconstruction performance. Figure˜8 shows the reconstruction error for varying number of latent vectors $L$. Even with substantially fewer latent vectors than entities, the model achieves good reconstruction performance. This gap suggests that the performance gains from increasing $L$ are not due to improved reconstruction, but rather to the model’s ability to leverage the enlarged latent space representation more effectively.
@@ G.4 Identifier Pool Size
We evaluate the impact of identifier pool size using the MD17 dataset. This dataset contains at most 21 atoms, so, in general, an identifier pool of $|{\mathcal{I}}|=21$ would be sufficient. However, for the results shown in the main paper, we used $|{\mathcal{I}}|=32$. To investigate the impact of a larger identifier pool $\mathcal{I}$}, we conduct additional experiments by training multiple first-stage models with varying identifier pool sizes and report the reconstruction error measured by Euclidean distance. Figure˜7 shows that the reconstruction error increases with the size of the identifier pool, since a larger pool results in fewer updates to each entity embedding during training.
@@ G.5 Identifier Assignment
To assess the impact of different ID assignments on reconstruction performance, we run our model five times with different random ID assignments and measure the standard deviation across these assignments in terms of reconstruction error in Å. Results are shown in Figures˜7 and 8. The low standard deviation across different ID assignments demonstrates the robustness of our model with respect to random ID assignment.
G.6 Identifier Latent Utilization
We investigate how identifiers are utilized by our model, addressing three key questions: Does each identifier exhibit a unique query pattern? How do these patterns vary with latent space dimensionality? Do different attention heads learn distinct patterns?
We examined decoder attention weights in overparameterized ($N<L$) and compressed ($N\geq L$) settings using models with $L=16$ and $L=32$ latent vectors trained on MD17. Our analysis focuses on Aspirin molecules ($N=21$ atoms) across different conformations.
The results in Figs.˜9, 10, 11 and 12 reveal two main findings. First, identical atom identifiers produce consistent attention patterns across different conformations, with each identifier learning a unique addressing scheme distributed over latent vectors. This suggests that the identifiers function as "retrieval mechanisms" independent of stored content.
Second, attention head utilization adapts to the compression ratio. The overparameterized model ($L=32$) primarily uses a single head, acting like a hash table-like retrieval. The compressed model ($L=16$) utilizes both heads with distinct patterns, likely to mitigate identifier collisions in the limited latent space.
@@ G.6 Identifier Latent Utilization
Figure 9: Attention patterns on the MD17 dataset for four random samples of the validation dataset, using a model trained with 16 latent vectors. Each row corresponds to a single sample and displays the attention patterns across individual attention heads. Each sample employs a distinct identifier assignment.
Figure 10: Attention patterns on the MD17 dataset for four random samples of the validation dataset, using a model trained with 16 latent vectors. Each row corresponds to a single sample and displays the attention patterns across individual attention heads. Each sample employs the same identifier assignment.
Figure 11: Attention patterns on the MD17 dataset for four random samples of the validation dataset, using a model trained with 32 latent vectors. Each row corresponds to a single sample and displays the attention patterns across individual attention heads. Each sample employs a distinct identifier assignment.
Figure 12: Attention patterns on the MD17 dataset for four random samples of the validation dataset, using a model trained with 16 latent vectors. Each row corresponds to a single sample and displays the attention patterns across individual attention heads. Each sample employs the same identifier assignment.